The cost of renting a banquet hall is given by the linear equation y=25x+100. Write this equation in standard form.
And
Brianna sells computers and earns base salary of $300 per week and $50 for every computer sold. Her salary is modelled by the equation y=50x+300. Write this linear equation in standard form.

Found 2 solutions by addingup, ankor@dixie-net.com:
Answer by addingup(3677)

-25x + y = 100
.
you do the other, the standard form of an equation is Ax+By = C where A,B, and C are integers.

Answer by ankor@dixie-net.com(22740)

The cost of renting a banquet hall is given by the linear equation y=25x+100. Write this equation in standard form.
y = 25x + 100
-25x + y = 100
And
Brianna sells computers and earns base salary of $300 per week and $50 for every computer sold.
Her salary is modelled by the equation y=50x+300.
Write this linear equation in standard form.
-50x + y = 300